maximum or minimum of f(x, y) = x3+ y3– 63(x + y) + 12xy



1
Expert's answer
2022-05-04T17:27:38-0400

The two variables function is given by

"F(x,y) = x^3 + y^3 - 63(x+y) + 12xy"


The first order partial derivatives of this function are

"F_x = 3x^2 - 63 + 12y \, \, and \,\, F_y = 3y^2 - 63 + 12x"


The second order partial derivatives of this function are

"F_{xx} = 6x \,\,\, , \,\,\, F_{yy} = 6y \,\,\, and \,\,\, F_{xy} = 12"


To find critical points of this function, we have to set "F_x = 0 \,\,\, and \,\,\, F_y = 0"

Hence:

"F_x = 0 \\
\Rightarrow 3x^2 - 63 + 12y = 0 \\
\Rightarrow 12y = 63 - 3x^2"

"\Rightarrow 4y = 21 - x^2 \\
\Rightarrow y = \frac{21 - x^2}{4}" (Equation 1)


And:

"F_y = 0 \\
\Rightarrow 3y^2 - 63 + 12x = 0 \\
\Rightarrow y^2 + 4x - 21 = 0"

"\Rightarrow \frac{(21 - x^2)^2}{16} + 4x - 21 = 0 \hspace{1 cm} \left[ \because y = \frac{21 - x^2}{4} \right]"

"\Rightarrow 441 - 42x^2 + x^4 + 64x - 336 = 0 \\
\Rightarrow x^4 - 42x^2 + 64x + 105 = 0 \\"

Using a graphing calculator, we get the solutions of this polynomial equation and which are:

"x = -7 \,\,\, , \,\,\, x = -1 \,\,\, , \,\,\, x = 3 \,\,\, and \,\,\, x = 5"


Inserting "x = -7" into the (Equation 1), we have:

"y = \frac{21 - (-7)^2}{4} \\
\Rightarrow y = -7"

Inserting "x = -1" into the (Equation 1), we have:

"y = \frac{21 - (-1)^2}{4} \\ 
\Rightarrow y = 5"

Inserting "x = 3" into the (Equation 1), we have:

"y = \frac{21 - 3^2}{4} \\ 
\Rightarrow y = 3"

Inserting "x = 5" into the (Equation 1), we have:

"y = \frac{21 - 5^2}{4} \\
 \Rightarrow y = -1"


Hence the critical points of this function are

"(-7, -7) \,\,\, , \,\,\, (-1,5) \,\,\, , \,\,\, (3,3) \,\,\, and \,\,\ (5, -1)"


Now,

to classify the critical points and obtain the relative extrema values of this function, we will use the following value of  "D"

"D = F_{xx} F_{yy} - (F_{xy})^2 \\
 \,\,\,\,\,\, = (6x)(6y) -(12)^2 \\
 \,\,\,\,\,\, = 36xy - 144"


At "(-7,-7)":

"D = 36(-7)(-7) - 144 \\
\,\,\,\,\,\,\, = 1620 > 0"

and:

"F_{xx} (-7,-7) \,= 6(-7) \\
\hspace{1 cm} \hspace{1 cm} = -42 < 0"

At "(-1,5)":

"D = 36(-1)(5) - 144 \\
 \,\,\,\,\,\,\, = -324 < 0"

At "(3,3)":

"D = 36(3)(3) - 144 \\
\,\,\,\,\,\,\, = 180> 0"

"F_{xx} (3,3) \,= 6(x) \\
\hspace{1 cm} \hspace{1 cm} = 18 > 0"

At "(5,-1)":

"D = 36(5)(-1) - 144 \\
 \,\,\,\,\,\,\, = -324 < 0"

So we can say that:

The saddle points of this function are

"(5,-1) \,\,\,\, and \,\,\,\, (-1,5)"

The relative maximum of this function exists at "(-7,-7)"

and

the relative minimum of this function exists at "(3,3)"


The relative maximum value of the function is

"F(-7,-7) = (-7)^3 + (-7)^3 - 63(-7-7)" "+ 12(-7)(-7)"

"\hspace{1 cm} \,\,\,\,\,\,\,\,\,\,\,\,\, = 784"

The relative minimum value of the function is

"F(3,3) = 3^3 + 3^3 - 63(3+3) + 12(3)(3)"

"\hspace{1cm} \,\,= -216"